OpenStudy (anonymous):

TWO TIMES THE BASE OF A PARALLELOGRAM IS EQUAL TO 3 TIMES THE SIDE. THE PARALLELOGRAM HAS A PERIMETER OF 800. WHAT IS ITS BASE? AND SIDE?

OpenStudy (anonymous):

If b is the base and s the side, you're told,\[2b=3s\]You also know that \[2b+2s=800\]You have two linearly independent equations in two unknowns - you can solve this system. Since the first equation gives 2b=3s, you can write for the second equation,\[2b+2s=3s+2s=800 \rightarrow 5s=800\]Therefore\[s=160\]Substitute this into the any one of the two original equations to get for the base,\[2b=3 \times 160 \rightarrow b=240\]

OpenStudy (anonymous):

CAN YOU PLEASE HELP WITH MY OTHER QUESTIONS

OpenStudy (anonymous):

Haven't other people answered them?

OpenStudy (anonymous):

NOT QUITE WELL

OpenStudy (anonymous):

Is it just the four questions in capitals?

OpenStudy (anonymous):

YES